8, Consider a pendulum bob of length 10cm hung from a ceiling. What should we do to the length in order to vary the period by 2?


1
Expert's answer
2021-09-16T05:14:01-0400

We know that, "T=2\pi \sqrt{\frac{L}{g}}"

Here, "L=10 \ cm" ", g=9.8 m\/s^2"

So, "T=2\pi\sqrt{\frac{10}{9.8}}=6.347\ s"

If "T" is varying by 2 then new length "(L_1)" will be calculated by

"T\pm2=2\pi \sqrt{\frac{L_1}{g}}"

"\Rightarrow 6.347\pm2=2\pi \sqrt{\frac{L_1}{9.8}}"

For "+" ve sign

"6.347+2=2\pi \sqrt{\frac{L_1}{9.8}}
\\\Rightarrow L_1=17.29\ m"

For "-" ve sign

"6.347-2=2\pi \sqrt{\frac{L_1}{9.8}}
\\\Rightarrow L_1=4.69\ m"